Obtain the partial differential equation by eliminating the arbitrary constant from the relation

x^2/a^2 +y^2/b^2 +u^2/c^2 =1


1
Expert's answer
2021-09-27T16:10:18-0400

"\frac{x^2}{a^2} + \frac{y^2}{b^2} + \frac{u^2}{c^2} = 1"

Differentiating w.r.t. x

"\frac{2x}{a^2} + \frac{2u}{c^2}\frac{du}{dx} = 0 \;\;\;(1)"

Again differentiating (1) w.r.t. x

"\frac{2}{a^2} + \frac{2ux}{c^2} \frac{d^2u}{dx^2} + \frac{2}{c^2} (\frac{du}{dx})^2 = 0"

Multiplying by x

"\frac{2x}{c} + \frac{2ux}{c^2} \frac{d^2u}{dx^2} + \frac{2}{c^2} (\frac{du}{dx})^2x = 0"

From (1):

"\frac{2x}{a^2} = \frac{-2u}{c^2} \frac{du}{dx} \\

\frac{-2u}{c^2} \frac{du}{dx} + \frac{2ux}{c^2} \frac{d^2u}{dx^2} + \frac{2x}{c^2} (\frac{du}{dx})^2 = 0"

Put

"\frac{du}{dx} = p \\

\frac{d^2u}{dx^2} = s \\

-up + uxs + xp^2 =0"